how can i solve y=3x+2 if the domain is (-2 , -1 , 0 , 2 , 5 )
Found 2 solutions by venugopalramana, sukusu:
Answer by venugopalramana(3286)

how can i solve y=3x+2 if the domain is (-2 , -1 , 0 , 2 , 5 )
domain means the values of x
range is the corresponding values of y
say for x=-2,we get y=3*-2+2=-6+2=-4
similarly find others and tabulate as below
x......-2.....-1......0........2.........5
y......-4.....-1......2........8.........17
so range is (-4,-1,2,8,17)...it is better represented as (-2,-4),(-1,-1),0,2),(2,8),(5,17)

Answer by sukusu(21)

replace x with the given numbers and solve the prolem
fro example lets take 2
y=3(2)+2
y=6+2
y=8
see if you can do the rest with the other domains